OpenStudy (tia_sweetheart17):

solve the solution: x^2-6x=40

OpenStudy (anonymous):

easiest to complete the square for this one \[(x-3)^2=40+3^2=40+9=49\]

OpenStudy (anonymous):

then \[x-3=7\] or \[x-3=-7\] solve those for your two answers